Solve the following initial value problem

x'(t) + y(t) = 2

y'(t) - x(t) = = δ(t − π)

x(0) = 0, y(0) = 1.
Answer #1

Consider a system of equations x+y= 2.-x+ y=0(t-1).x(0)=0,y(0)=1 Applying Laplace transform on both sides, we get (sz {x{t}}1+ ਬ $+-+se L{y()} = -+ ਨ+ 2 +1s (s +1) s** (2)=rਜ ਤੇ ਨੂੰ y(1) =-cost+2+1 (1-1)cos (1-1)